You have the following information for Hippea Chippea Company: current stock price = $81; required rate of return = 11.1%; growth rate in dividends = 3%. What was the most recent dividend per share paid on the stock?
Answer #1

According to the Dividend growth model,

The Dividend Discount Model Still Going Strong - DivHut

We need to calculate D0.

D.(1+9) Po = (r-9) Por-g)_$81* (11.1% -3%) = $6.37 (1 + 9) (1+3%) Do =

Hence, the most recent dividend per share paid on the stock is $6.37